calculate the sum of the interior angles of a polygon with 7sides,9sides,20sides,
Answer by Earlsdon(6103)

I'll show you how to do one of them and you can do the others.
You can derive the formula yourself by doing the following:
1) In any polygon, from one vertex (corner), draw all the diagonals to the other vertices.
2) Now count the number of triangles you have made inside the polygon. This number will always be 2 less than the number of sides in the polygon.
Each triangle has a total of 180 degrees as the sum of its interior angles.
3) Multiply the number of triangles you made by 180 degrees and this will be the sum of the interior angles of the polygon.
The formula is:
(n-2)180 degrees. Where n is the number of sides of the polygon.
Let's do the 20-gon, (icosagon?)
The number of sides is 20, so n = 20.
(n-2)180 degrees = (20 - 2)180 = (18)180 = 3240 degrees.
You can do the others using the same formula.